OpenStudy (anonymous):

the sum of the reciprocal of two numbers is 3/8. find the other number, if one is 4 greater than the other.

OpenStudy (anonymous):

1/x +1/y=3/8 x+4=y 1/x+1/(x+4)=3/8 x+4+x=3/8(x)(x+4) 16x+32=3x^2+12x 3x^2-4x-32=0 x= 4 or x= -8/3 y=8 y=4/3